OpenStudy (anonymous):

HELP ! if the median height of 12 students is 5’6 how many students must be exactly 5’6 A-0 B-2 C-6 D-12

OpenStudy (amistre64):

does the middle of an even set have to be a part of the set?

OpenStudy (amistre64):

does an average of 2 numbers, have to be equal to either number?

OpenStudy (anonymous):

no?

OpenStudy (amistre64):

thats correct

OpenStudy (amistre64):

the median of an even set of number will be the average of the 2 numbers in the middle

OpenStudy (anonymous):

I am having trouble with how I figure out the answer. I think it would be 0, because the it is the median,

OpenStudy (amistre64):

the answer is zero. none of them have to be exactly 5'6

OpenStudy (amistre64):

the average of 5'5 and 5'7, with any luck ... is 5'6, which is not a member of the set to start with

OpenStudy (anonymous):

Oh okay, thank you for your help! Could you help me with one more?

OpenStudy (amistre64):

maybe .. my memory can be rather sporatic

OpenStudy (anonymous):

You are keeping track of your average daily caloric intake in order to begin reducing the number of unnecessary calories in your daily diet. Your average for the past month (30 days) 2as 1926 calories a day. You just returned from a 5 day vacation and you would like to recalculate your average. On vacation, you consumed on average of 2500 calories a day. What is your new average? A. 574 B. 1944 C. 2008 D. 2213

OpenStudy (amistre64):

how do we caluclate an average?

OpenStudy (anonymous):

You add up all of the numbers, and then divide it by how many numbers there are

OpenStudy (amistre64):

good, so:\[\frac{A}{30}=1926\] \[A=1926(30)\] adding in the next 5 days ....\[\frac{A+5(2500)}{35}=New_{avg}\]

OpenStudy (anonymous):

So would I add 30+5(2500)/35?

OpenStudy (amistre64):

not quite. the sum of the first 30 days is: 1926(30) the sum of the next 5 days is: 5(2500) add them up for the total sum of 35 days

OpenStudy (anonymous):

so I would do 1926x30? and then 5(2500)

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

I get and 57,780 12,500

OpenStudy (amistre64):

i get: 1926(30) + 5(2500) = 70280 70250/35 is one of the options

OpenStudy (anonymous):

Would it be c?

OpenStudy (amistre64):

i get C. with any luck its also correct :)

OpenStudy (anonymous):

Awesome! thank you for your help

OpenStudy (amistre64):

youre welcome, and good luck :)

OpenStudy (anonymous):

Thank you!